3.1. Convert the contingency table to show the conditional distribution of cigarette use. Show your work. Hint: Find P(cigarette=yes|alcohol=yes) etc.

Alcohol Use Cigarette Use
Yes No
Yes 1200 400
No 300 200
Answer #1

joint distribution is :

Yes No total
yes 1200 400 1600
No 300 200 500

dividing each value in joint distribution with row total, below is conditon distribution of cigrette use:

Yes No
yes 1200/1600=0.75 400/1600=0.25
No 300/500=0.6 200/500=0.4

from above:  P(cigarette=yes|alcohol=yes) =0.75

P(cigarette=no|alcohol=yes) =0.25

P(cigarette=yes|alcohol=no) =0.60

P(cigarette=no|alcohol=no) =0.4